# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ III.6 High-performance real-time decoding platform
Challenges and requirements. The high speed of superconducting processors, a great advantage for utility-scale applications, requires well-engineered control and decoder architecture, both on the software and hardware levels. A key technical challenge is that decoding simultaneously requires peta-scale computation and low latency for real-time feed-forward. Furthermore at stage, it isn’t known yet what algorithms are most effective at decoding, therefore there is a systems engineering trade-off between on performance and flexibility.
@@
Performing universal fault-tolerant quantum computation using QEC mandates feedforward-based implementation of certain quantum gates (e.g. $T$ gates) with low latency Kurman et al. (2023). In these implementations, a conditional operation is applied based on the result of a logical measurement as well as the decoding of syndromes of many previous QEC cycles. The classical feed-forward latency is measured from the physical execution of the logical measurement until the controller executes a conditional gate (L0 and L1 in Figure 25(c).
For efficient execution of fault-tolerant feed-forward gates, the decoder needs to be ready in time for the conditional gate execution. We note that on average, $d$ (distance) cycles are allowed for the decoder result for multiple reasons. First, when the conditional gate is followed by gates that commute, it may be deferred after the gates that do not depend on the decoding result. Second, the gates that follow the conditional gate may require synchronization with other surfaces, allowing to defer the conditional gate without impacting the circuit. In addition, we note that sporadic delays caused by the decoder have a small impact on the overall performance as long as on average the decoder result is ready on time, as shown in Figure 25(b). Therefore we target an end-to-end average decoding latency shorter than $d$ QEC cycles, which implies a target latency of approximately 10 us. To meet these latency targets and implement QEC decoding efficiently, it is essential to optimize the performance of both the control-decoder communication channel and decoding task. For the controller-decoder channel, throughput must exceed the data generation rate. The controller should locally perform state discrimination including optional soft readout indicators, encoding each qubit state with a minimal bit representation. With 20K qubits, a 4-bit state representation per qubit, and a QEC cycle time of 550 ns, a minimum net bandwidth of 150 Gb/s is required. Additionally, data sent from the decoder to the controller must be efficiently compacted to communicate only the necessary logical instruction. In addition, the overall latency should be minimal, including readout state discrimination, data aggregation from multiple controllers and transmission to and from the decoder.
@@
The decoding processes, which include multiple concurrent decoding processes, should be designed to minimize latency. Scalable hardware is required, as decoding for circuits with 10K–100K qubits demands extensive computational capacity. Some decoding algorithms, such as Fusion Blossom, may exhibit variability in the decoding time dependent on the error pattern. The QEC implementation should be designed to accommodate this variability. For the decoder not to limit the performance, the average throughput of the decoder should exceed the syndrome generation rate. In addition, the decoder average latency including the roundtrip communication time should be shorter than the d QEC cycles.
@@
Figure 25: Example of non-Clifford computation with surface codes. (a) An example of a logical circuit containing two non-Clifford gates. (b) The fault-tolerant logical circuit that implements the circuit in (a) with surface codes with only a single ancillary surface. The dashed square denotes the feed-forward conditional logical gates that verify that the planned circuit is executed. (c) The space–time view of the circuit in (b) with surface codes. Each colour denotes a separate decoding task, chosen to end each task with a logical measurement. The decoding outcome of the lattice surgery between a magic state surface and the computation surface determines a feed-forward circuit, which delays the circuit by if the feed-forward latency ($L$) is larger than a threshold latency ($L_{th}$).
Real-time decoding architecture. To address the FTQC requirements, a proposed architecture, illustrated in 12 is designed to allow a high-performance decoding platform along with low latency communication between the quantum control clusters and the decoding platform. The control clusters control one or more quantum surfaces, operating in synchronization to execute QEC cycles based on the state of each surface. In addition, the control clusters are responsible to benchmark the qubits performance and maintain qubits calibration. The acceleration servers provide a high-performance platform for the execution of decoder instances and the logical circuit orchestration. They are based on CPU/GPU processors with direct data transfer capabilities to and from the control system. We note that GPUs are beneficial for real-time decoding of QECCs thanks to their massive parallelism and their high-bandwidth/low-latency interfaces. In addition, the servers may incorporate specialized ASIC or FPGA acceleration cards and dedicated hardware offload capabilities. In large-scale systems, multiple acceleration servers may operate in parallel, running multiple decoder instances. The control clusters and accelerator servers are connected by a low-latency network. The network facilitates the aggregation of readout data from the control clusters and distribution of the data to the appropriate decoders. To meet the end-to-end latency requirements, the total time for data aggregation, roundtrip communication and decoding should be in the order of 10 us, which require the design of a specialized communication protocol for QEC.
@@
Real-time decoding on a DGX Quantum platform. DGX Quantum provides a tight integration of QM’s OPX1000 controller with Nvidia’s Grace Hopper (GH) superchip and offers an effective platform for FTQC, supporting both logical circuit orchestration and decoding processes. The close integration of CPU and GPU resources enables real-time, parallel execution of various decoding algorithms, including deep-learning-based distributed decoders. The system is connected to the low-latency QEC network and transfers data from the control system to the high-performance CPU-GPU and vice-versa over a PCIe interface. A round-trip feed-forward latency (not including the decoding task) from measurement to the decoder and back to conditional gate control, has been benchmarked at less than 3.8 µs.
The DGX Quantum platform is designed to be connected to a hierarchical, scalable QEC network for data aggregation and distribution, ensuring low latency across systems with 10K qubits and beyond. The standard, modern PCIe interface provides 100s of Gb/s of bandwidth and supports also communication with connected ASIC or FPGA acceleration cards. As scaling extends to systems with 10K–100K qubits, a dedicated, optimized interface connecting the QEC network directly to the decoders may be desired, to further reduce the latency.
@@
DGX Quantum leverages the extensive parallel processing capabilities, large memory, and high memory bandwidth of the GPU, providing a robust platform for QEC decoding and runtime execution. Moreover, as a software based solution, DGX Quantum provides a platform for flexibility and rapid development that are desired in the early stages of FTQC.
@@
Preliminary evaluation of a software-based implementation of Fusion Blossom algorithm with batch decoding on a DGX Quantum server demonstrated that a serial implementation could sustain the necessary decoding throughput for a distance d=11, with a basic error model with error probability $P_{error}=1\texttimes 10^{-3}$. In the next steps, we plan to implement Fusion Blossom in stream mode, in addition to leveraging parallel processing and utilizing the large memory capacity for potential caching and optimization for common error patterns.
@@ III.7 Distributed FTQC across multiple dilution refrigerators
A fault-tolerant quantum computer with 1M+ physical qubits may require multiple dilution refrigerators (DR) with quantum interconnects between DRs. The inter-DR and intra-DR architecture of the computer is discussed further in Section III.7. The assembler prioritizes inter-DR lattice surgeries (involving less than 120k qubits) over multi-fridge surgeries, as logical teleportation of states between different DRs is much slower and of lower fidelity than intra-DR operations. Multi-DR surgeries involving code distance $d$ will require at least $d$ optical interconnects between nearest-neighbour DRs, which is a demanding requirement Lee et al. (2020a); Youssefi et al. (2021); Lecocq et al. (2021); Awschalom et al. (2021); Delaney et al. (2022).
Assembling large FTQC programs among multiple DRs. Embedding FTQC architectures across multiple DRs involves solving a complex embedding problem to determine the connectivities between DRs, teleporation sites adjacent to the interconnects, and appropriate areas across the multi-DR system for the core processor and the MSF zones of required code distances. The embedding prioritizes intra-DR connectivity to ensure the robustness of FTQC protocols against noise introduced by the imperfect interconnects. This is also an important consideration even within individual DRs when lattice surgeries span across the edge couplers of the QPU (i.e., the weaker capacitive coupling between the 20k qubit wafers).
@@
Figure 26: Example of an embedded multi-DR architecture for executing a quantum circuit associated with electronic-structure quantum simulation based on qubitization of the $p$-benzyne molecule for an active space involving six molecular orbitals (see Section IV). Our estimates indicate that this circuit requires at least 11 distillation units in the MSF with a code distance of 21 to ensure a continuous supply of magic states to the core processor, which requires 341 data qubits with a code distance of 25. The architecture shown consists of 15 DRs, each containing 120,000 physical qubits. Four DRs are configured with three distillation units each, 10 DRs accommodate 33 data qubits each, and the remaining DR includes additional data qubits along with dedicated zones for magic state growth and storage. Multi-qubit lattice surgery is performed using the quantum bus, while magic states are transferred between DRs using as many optical interconnects as the code distances involved.
@@
In Figs. 28(a) and (b), we show that using weaker CNOTs of infidelity $p_{\text{link}}=0.01$ and fixing all other coupler infidelities to the baseline value $p_{CX}=0.003$ does not significantly affect the teleportation fidelity even for 4 cuts within the bus. However, much weaker interconnects can be problematic as shown in Figs. 28(c) and (d) where we observe a thresholding behavior at about $p_{\text{link}}\approx 0.06$.
We conclude that distributed surface code architectures across multiple dilution refrigerators can tolerate two-qubit errors on the order of $1\%$ arising from noisy optical interconnects between the DRs. However, the logical performance rapidly deteriorates as these errors become significantly worse, and for error rates beyond $5\%$, achieving fault tolerance becomes problematic.
Figure 28: (a) Logical infidelity as a function of code distance for teleportation of $|+_{L}\rangle$ and $|0_{L}\rangle$ states with and without cuts for corresponding values of bus width $b=3d$ and temporal code distance $r_{m}=3d$. (b) Exponential fits for $P_{0}$ and $P_{+}$ for $n_{\text{cuts}}=4$. The obtained $\mu$ and $\Lambda$ values are close to that of Figure 24(b). Logical infidelity of teleporting $|0_{L}\rangle$ (c) and $|+_{L}\rangle$ states (d) as a function of the infidelity of the weak $CX$ gates in the cuts.
@@ IV.1 Quantum computation of electronic spectra as a representative high-utility application
One of the most promising applications of quantum computing is solving quantum chemistry problems. An important representative computational task in quantum chemistry is estimating ground-state energies of molecules. While this task is classically tractable for small molecules using advanced classical algorithms developed in the field of traditional quantum chemistry Szabo and Ostlund (1996), electronic-structure simulations of large molecules are widely considered to be intractable for classical computers. Here, for the purpose of demonstrating the practicality of solving such problems on a quantum computer, we present QRE studies of electronic-structure quantum simulations for two molecules of high practical interest. The first molecule analyzed is the biradical para-benzyne molecule ($p$-benzyne), which has the molecular formula $\mbox{C}_{6}\mbox{H}_{4}$. Its energetically lowest configuration is formed by a singlet biradical. Among other applications, its reactivity has the potential to play an important role in the design of enediyne drugs with high antitumour or anticancer properties Crawford et al. (2001); Kraka and Cremer (2000). The second molecule analyzed is the iron-molybdenum cofactor (FeMoco) of nitrogenase, which acts as a crucial catalyst in the process of biological nitrogen fixation. This molecule, as well as many others, have been used as representative targets for future quantum simulators in several recent works Reiher et al. (2017); Motta et al. (2021a); Li et al. (2019); Lee et al. (2021); Otten et al. (2023).
@@
Our studies rely on electronic-structure simulations in the second quantization framework of quantum theory. Numerous software tools exist to derive the second-quantized Hamiltonian from a molecule’s specifications that fully characterize the quantum system. Basic molecular specifications include the types of atoms that constitute the molecule and the molecule’s geometry (typically summarized in an xyz file), total charge, and the total spin. In addition, a basis set $\{\phi_{\alpha}(\bm{x})\}$ must be selected to represent the fermionic orbitals, which in the language of second quantization are occupied or unoccupied, represented by occupation number states and fermionic creation and annihilation operators ($\hat{a}_{\alpha}$ and $\hat{a}^{\dagger}_{\alpha}$) acting upon them. Furthermore, to reduce the computational cost, a common approach is to restrict computations to a reasonably chosen active space involving only a subset of the chosen orbital basis set. Finally, the model Hamiltonian associated with an active space is translated from the second quantization framework to a framework suitable for the quantum circuit model. This fermion-to-qubit mapping is typically accomplished via either the Jordan–Wigner Jordan and Wigner (1993) or the Bravyi–Kitaev Bravyi and Kitaev (2002) transformations. To derive the model Hamiltonians for the various active spaces associated with $p$-benzyne, we used Tangelo Senicourt et al. (2022), an open source Python software package for end-to-end chemistry workflows. For the FeMoco molecule, we used the FCIDUMP file provided as part of the data and code repository Lee et al. (2020b) of Ref. Lee et al. (2021).
@@
The standard QPE algorithm Nielsen and Chuang (2011) samples in the eigenbasis of the molecular Hamiltonian $H$ by measuring the phase that is accumulated on an initial input quantum state through multiple controlled executions of the time-evolution operator $\exp(-iHt)$. Its most resource-intensive part consists in implementing the unitary $\exp(-iHt)$ by a quantum circuit, along with repeating this circuit a number of times that scales as $\mathcal{O}(1/\epsilon)$ for an allowable target error $\epsilon$ in phase estimation. An alternative approach to sampling the spectrum of the molecular Hamiltonian $H$ via phase estimation is based on the framework of qubitization Low and Chuang (2019). Indeed, most of the recent QRE studies on electronic-structure quantum simulations rely on the qubitization framework (see, e.g., Berry et al. (2019); von Burg et al. (2021a); Lee et al. (2021); Goings et al. (2022); Beverland et al. (2022); Otten et al. (2024); Bellonzi et al. (2024); Watts et al. (2024); Nguyen et al. (2024)). This approach uses a new operation called qubiterate that is akin to the quantum walk operator $e^{i\arccos(H/\lambda)}$ (where $\lambda$ is typically the sum of the absolute values of the weightings in the molecular Hamiltonian). Since the qubiterate’s eigenvalue spectrum can be obtained from that of the unitary $\exp(-iHt)$ via an $\arccos$ transformation, the former can be used in QPE in place of the latter, as proposed in Refs. Berry et al. (2018); Poulin et al. (2018). An advantage of this approach is that steps of a quantum walk can be implemented exactly, assuming access to arbitrary single-qubit rotations, in contrast to all approaches that are based on Hamiltonian simulation of the time-evolution operator $\exp(-iHt)$, which can only be approximated.
@@
Moreover, aiming to reduce algorithmic complexity, the majority of recent literature on electronic-structure quantum simulations and the associated resource requirements has focused on combining the technique of qubitization applied to molecular systems with various tensor factorization techniques for the Coulomb operator. State-of-the-art algorithms of this type include the single low-rank factorization algorithm of Berry et al. Berry et al. (2019), the double low-rank factorization algorithm of von Burg et al. von Burg et al. (2021a), and the tensor hypercontraction algorithm of Lee et al. Lee et al. (2021), resulting in a continual improvement of the $T$ gate or Toffoli gate complexity from $\mathcal{O}\left(N^{5}/\epsilon^{3/2}\right)$ for the Trotter-based approach to $\mathcal{O}\left(N\lambda/\epsilon\right)$, where $\lambda$ is the 1-norm of Hamiltonian coefficients which typically has a scaling between $\mathcal{O}\left(N\right)$ and $\mathcal{O}\left(N^{3}\right)$.
IV.2 Quantum resource estimation studies for $p$-benzyne and FeMoco

### v2

@@ III.6 High-performance real-time decoding platform
Challenges and requirements. The high speed of superconducting processors, a great advantage for utility-scale applications, requires well-engineered control and decoder architecture, both on the software and hardware levels. A key technical challenge is that decoding simultaneously requires peta-scale computation and low latency for real-time feed-forward. Furthermore, at stage, it is not known yet what algorithms are most effective at decoding; therefore, there is a systems engineering trade-off between performance and flexibility.
@@
Performing universal fault-tolerant quantum computation using QEC mandates feedforward-based implementation of certain quantum gates (e.g., $T$ gates) with low latency Kurman et al. (2024a). In these implementations, a conditional operation is applied based on the result of a logical measurement as well as the decoding of syndromes of many previous QEC cycles. The classical feed-forward latency is measured from the physical execution of the logical measurement until the controller executes a conditional gate ($L^{0}$ and $L^{1}$ in Figure 26(c)).
[…]
@@
The decoding process, which includes multiple concurrent decoding tasks Kurman et al. (2024b), should be designed to minimize the result latency. Scalable hardware is required, as decoding for circuits with 10k–100k qubits demands extensive computational capacity. Some decoding algorithms, such as Fusion Blossom, may exhibit variability in the decoding time dependent on the error pattern. The QEC implementation should be designed to accommodate this variability. For the decoder to not limit performance, the average decoder throughput should exceed the syndrome generation rate. In addition, the average decoder latency, including the roundtrip communication time, should be shorter than the $d$ QEC cycles.
@@
Figure 26: Example of non-Clifford computation with surface codes adapted from Ref. Kurman et al. (2024a). (a) Example of a logic circuit containing two non-Clifford gates. (b) Fault-tolerant logic circuit that implements the circuit in (a) with surface codes with only a single ancillary surface. The dashed square denotes the feed-forward conditional logical gates that verify that the planned circuit is executed. (c) Space–time view of the circuit in (b) with surface codes. Each colour denotes a separate decoding task, chosen to end each task with a logical measurement. The decoding outcome of the lattice surgery between a magic state surface and the computation surface determines a feed-forward circuit, which delays the circuit by if the feed-forward latency ($L$) is larger than a threshold latency ($L_{\text{th}}$).
[…]
@@
Real-time decoding on a DGX Quantum platform. DGX Quantum provides tight integration of QM’s OPX1000 controller with Nvidia’s Grace Hopper (GH) superchip and offers an effective platform for FTQC, supporting both logic circuit orchestration and decoding processes (see Appendix F for details). The close integration of CPU and GPU resources enables real-time, parallel execution of various decoding algorithms, including deep-learning-based distributed decoders. The system is connected to the low-latency QEC network and transfers data from the control system to the high-performance CPU–GPU and vice versa over a PCIe interface. A round-trip feed-forward latency (not including the decoding task) from measurement to the decoder and back to conditional gate control, has been benchmarked at less than 3.8 µs.
The DGX Quantum platform is designed to be connected to a hierarchical, scalable QEC network for data aggregation and distribution, ensuring low latency across systems with 10K qubits and beyond. The future DGX Quantum platform, based on the Blackwell architecture B200 GPUs, utilizes PCIe 6.0, which provides up to 128 GB/s bidirectional bandwidth in an x16 configuration. The PCIe 6.0 specification incorporates Forward Error Correction (FEC) to maintain data integrity and limit additional latency to under two nanoseconds, ensuring high-speed data transfers with low latency. As scaling extends to systems with 10K–100K qubits, a dedicated, optimized interface connecting the QEC network directly to the decoders may be desired to further reduce the latency. B200 GPU achieves up to 18 petaflops (PFLOPS) for sparse operations. This is made possible through advanced sparsity techniques and architectural enhancements, doubling the effective performance in scenarios where sparsity can be exploited effectively, which is often the case for quantum circuit simulations given the sparsity of data.
@@
DGX Quantum leverages the extensive parallel processing capabilities, large memory, and high memory bandwidth of the GPU, providing a robust platform for QEC decoding and runtime execution. Moreover, as a software-based solution, DGX Quantum provides a platform for flexibility and rapid development that is desired in the early stages of FTQC.
@@
Preliminary evaluation of a software-based implementation of the Fusion Blossom algorithm with batch decoding on a DGX Quantum server demonstrated that a serial implementation could sustain the necessary decoding throughput for a distance d=11, with a basic error model with error probability $P_{error}=1\texttimes 10^{-3}$. In the next steps, we plan to implement Fusion Blossom in stream mode, in addition to leveraging parallel processing and utilizing the large memory capacity for potential caching and optimization for common error patterns.
The DGX Quantum platform is particularly effective for QEC schemes that benefit from local decoding, as these can be efficiently accelerated by GPUs. Local decoders, such as Steiner tree-based decoders, are prime examples of algorithms that perform well in GPU-accelerated environments due to their parallelizable structure. In addition to standard local decoding, there are more-complex classes of quantum codes, such as entanglement-assisted quantum error-correcting codes (EAQECC) and entanglement-assisted quantum low-density parity-check (EA-QLDPC) codes, which offer enhanced error-correction capabilities by leveraging entanglement Panteleev and Kalachev (2022); Swaroop et al. (2024). Some of these codes, however, have varying degrees of GPU acceleration potential, depending on the complexity and structure of their decoding algorithms. Some EA-QLDPC codes, particularly those with irregular or non-local error syndromes, require more-sophisticated scheduling but can still benefit from GPU acceleration with optimized parallelization strategies. Recent research efforts have focused on discovering new classes of quantum codes that not only offer high error-correction rates but are also capable of full utilization of GPUs.
AI-assisted decoders and quantum algorithms can also significantly improve real-time decoding performance by leveraging rich AI infrastructure and GPU acceleration NVIDIA Corporation (2025a). A great example of AI’s application to QEC is Google’s AlphaQubit approach, which leverages machine learning to enhance the decoding process for QEC codes Pan (2024). AlphaQubit uses reinforcement learning and neural networks to identify optimal error-correction strategies by learning from a large amount of synthetic and experimental data.
@@ III.7 Distributed FTQC across multiple dilution refrigerators
A fault-tolerant quantum computer with 1M+ physical qubits may require multiple dilution refrigerators (DR) with quantum interconnects between the DRs. The inter-DR and intra-DR architecture of the computer is discussed further in Section III.7. The assembler prioritizes inter-DR lattice surgeries (involving less than 120k qubits) over multi-fridge surgeries, as logical teleportation of states between different DRs is much slower and of lower fidelity than intra-DR operations. Multi-DR surgeries involving code distance $d$ will require at least $d$ optical interconnects between nearest-neighbour DRs, which is a demanding requirement Lee et al. (2020a); Youssefi et al. (2021); Lecocq et al. (2021); Awschalom et al. (2021); Delaney et al. (2022).
Assembling large FTQC programs among multiple DRs. Embedding FTQC architectures across multiple DRs involves solving a complex embedding problem to determine the connectivities between DRs, teleporation sites adjacent to the interconnects, and appropriate areas across the multi-DR system for the core processor and the MSF zones of required code distances. The embedding prioritizes intra-DR connectivity to ensure the robustness of FTQC protocols against noise introduced by the imperfect interconnects. This remains an important consideration even within individual DRs when lattice surgeries span across the edge couplers of the QPU (i.e., the weaker capacitive coupling between the 20k qubit wafers).
@@
Figure 27: Example of an embedded multi-DR architecture for executing a quantum circuit associated with electronic-structure quantum simulation with a target error of 0.01 mHa based on qubitization of the $p$-benzyne molecule for an active space involving six molecular orbitals (see Section IV). Our estimates indicate that this circuit requires at least 11 distillation units in the MSF with a code distance of 21 to ensure a continuous supply of magic states to the core processor, which requires 341 data qubits with a code distance of 25. The architecture shown consists of 15 DRs, each containing 120,000 physical qubits. Four DRs are configured with three distillation units each, 10 DRs accommodate 33 data qubits each, and the remaining DR includes additional data qubits along with dedicated zones for magic state growth and storage. Multi-qubit lattice surgery is performed using the quantum bus, while magic states are transferred between DRs using as many optical interconnects as the code distances involved.
@@
In Figs. 29(a) and (b), we show that using weaker CNOT gates of infidelity $p_{\text{link}}=0.01$ and fixing all other coupler infidelities to the baseline value $p_{CX}=0.003$ does not significantly affect the teleportation fidelity even for four cuts within the bus. However, much weaker interconnects can be problematic as shown in Figs. 29(c) and (d) where we observe a thresholding behavior at about $p_{\text{link}}\approx 0.06$.
We conclude that distributed surface code architectures across multiple dilution refrigerators can tolerate two-qubit errors on the order of $1\%$ arising from noisy optical interconnects between the DRs. However, the logical performance rapidly deteriorates as these errors become significantly worse, and for error rates beyond $5\%$, achieving fault tolerance becomes problematic. Our numerical results are consistent with an analogous analysis performed for neutral-atom computers in Sinclair et al. (2024).
We note that the approach based on multi-DR lattice surgeries described in this section competes with the entanglement-distillation based quantum networking approach proposed in Ref. Sunami et al. (2024). In this approach, time-multiplexed remote physical Bell pairs are used to encode noisy logical Bell pairs by state injection, followed by entanglement distillation using logical operations on the code blocks to reach a desired error rate for the teleported state in the target QPU. However, the entanglement distillation units create a resource trade-off between the number of optical interconnects and the number of qubits used to complete the teleportation scheme in the target QPU. The analysis of such trade-offs is left for future resource estimations.
Figure 29: (a) Logical infidelity as a function of code distance for teleportation of $|+_{L}\rangle$ and $|0_{L}\rangle$ states with and without cuts for corresponding values of bus width $b=3d$ and temporal code distance $r_{m}=3d$. (b) Exponential fits for $P_{0}$ and $P_{+}$ for $n_{\text{cuts}}=4$. The obtained $\mu$ and $\Lambda$ values are close to that of Figure 25(b). (c)–(d) Logical infidelity of teleporting $|0_{L}\rangle$ states and $|+_{L}\rangle$ states, respectively, as a function of the infidelity of the weak $CX$ gates in the cuts.
@@ IV.1 Quantum computation of electronic spectra as a representative high-utility application
One of the most promising applications of quantum computing is sol quantum chemistry problems. An important representative computational task in quantum chemistry is estimating ground-state energies of molecules. While this task is classically tractable for small molecules using advanced classical algorithms developed in the field of traditional quantum chemistry Szabo and Ostlund (1996), electronic-structure simulations of large molecules are widely considered to be intractable for classical computers. Here, for the purpose of demonstrating the practicality of solving such problems on a quantum computer, we present QRE studies of electronic-structure quantum simulations for two molecules of high practical interest. The first molecule analyzed is the biradical para-benzyne molecule ($p$-benzyne), which has the molecular formula $\mbox{C}_{6}\mbox{H}_{4}$. Its energetically lowest configuration is formed by a singlet biradical. Among other applications, its reactivity has the potential to play an important role in the design of enediyne drugs with high antitumour or anticancer properties Crawford et al. (2001); Kraka and Cremer (2000). The second molecule analyzed is the iron-molybdenum cofactor (FeMoco) of nitrogenase, which acts as a crucial catalyst in the process of biological nitrogen fixation. This molecule, as well as many others, have been used as representative targets for future quantum simulators in several recent works Reiher et al. (2017); Motta et al. (2021a); Li et al. (2019); Lee et al. (2021); Otten et al. (2023).
@@
Our studies rely on electronic-structure simulations in the second quantization framework of quantum theory. Numerous software tools exist to derive the second-quantized Hamiltonian from a molecule’s specifications that fully characterize the quantum system. Basic molecular specifications include the types of atoms that constitute the molecule and the molecule’s geometry (typically summarized in an xyz file), total charge, and the total spin. In addition, a basis set $\{\phi_{\alpha}(\bm{x})\}$ must be selected to represent the fermionic orbitals, which in the language of second quantization are occupied or unoccupied, represented by occupation number states and fermionic creation and annihilation operators ($\hat{a}_{\alpha}$ and $\hat{a}^{\dagger}_{\alpha}$) acting upon them. Furthermore, to reduce the computational cost, a common approach is to restrict computations to a reasonably chosen active space involving only a subset of the chosen orbital basis set. The model Hamiltonian associated with an active space is translated from the second quantization framework to a framework suitable for the quantum circuit model. This fermion-to-qubit mapping is typically accomplished via either the Jordan–Wigner Jordan and Wigner (1993) or the Bravyi–Kitaev Bravyi and Kitaev (2002) transformations. To derive the model Hamiltonians for the various active spaces associated with $p$-benzyne, we used Tangelo Senicourt et al. (2022), an open source Python software package for end-to-end chemistry workflows. For the FeMoco molecule, we used the FCIDUMP file provided as part of the data and code repository Lee et al. (2020b) of Ref. Lee et al. (2021).
@@
The standard QPE algorithm Nielsen and Chuang (2011) samples in the eigenbasis of the molecular Hamiltonian $H$ by measuring the phase that is accumulated on an initial input quantum state through multiple controlled executions of the time-evolution operator $\exp(-iHt)$. Its most resource-intensive part consists in implementing the unitary $\exp(-iHt)$ by a quantum circuit, along with repeating this circuit a number of times that scales as $\mathcal{O}(1/\epsilon)$ for an allowable target error $\epsilon$ in phase estimation. An alternative approach to sampling the spectrum of the molecular Hamiltonian $H$ via phase estimation is based on the framework of qubitization Low and Chuang (2019). Indeed, most of the recent QRE studies on electronic-structure quantum simulations rely on the qubitization framework (see, e.g., Refs. Berry et al. (2019); von Burg et al. (2021a); Lee et al. (2021); Goings et al. (2022); Beverland et al. (2022); Otten et al. (2024a); Bellonzi et al. (2024a); Watts et al. (2024a); Nguyen et al. (2024a)). This approach uses a new operation called qubiterate that is akin to the quantum walk operator $e^{i\arccos(H/\lambda)}$ (where $\lambda$ is typically the sum of the absolute values of the weightings in the molecular Hamiltonian). Since the qubiterate’s eigenvalue spectrum can be obtained from that of the unitary $\exp(-iHt)$ via an $\arccos$ transformation, the former can be used in QPE in place of the latter, as proposed in Refs. Berry et al. (2018); Poulin et al. (2018). An advantage of this approach is that steps of a quantum walk can be implemented exactly, assuming access to arbitrary single-qubit rotations, in contrast to all approaches that are based on Hamiltonian simulation of the time-evolution operator $\exp(-iHt)$, which can only be approximated.
@@
Moreover, aiming to reduce algorithmic complexity, the majority of recent literature on electronic-structure quantum simulations and the associated resource requirements has focused on combining the technique of qubitization applied to molecular systems with various tensor factorization techniques for the Coulomb operator. State-of-the-art algorithms of this type include the single low-rank factorization algorithm of Berry et al. Berry et al. (2019), the double low-rank factorization algorithm of von Burg et al. von Burg et al. (2021a), and the tensor hypercontraction algorithm of Lee et al. Lee et al. (2021), resulting in a continual improvement of the $T$ gate or Toffoli gate complexity from $\mathcal{O}\left(N^{5}/\epsilon^{3/2}\right)$ for the Trotter-based approach to $\mathcal{O}\left(N\lambda/\epsilon\right)$, where $\lambda$ is the 1-norm of Hamiltonian coefficients which typically has a scaling between $\mathcal{O}\left(N\right)$ and $\mathcal{O}\left(N^{3}\right)$, and $\epsilon$ specifies the target error in ground-state energy estimation.
IV.2 Quantum resource estimation for $p$-benzyne and FeMoco